# E-TRIALS: Empowering Data-Driven Decisions to Enhance Computer-Based Learning Platforms

### v1

@@ Abstract
Computer-based learning platforms (CBLPs) have become a common medium in schools, transferring how students learn and interact with education content. A recent report by the Institute of Education Sciences revealed that 94% of public schools now provide a computer for every student. The high percentage of students having their own computers highlights the need for researchers to investigate how students learn in CBLPs and identify instructional designs that can enhance their performance. In this paper, we introduce E-TRIALS, a tool by ASSISTments designed to conduct experiments and analyze student performance. Using data from three experiments, we propose a generalizable approach for analyzing experimental data and modeling results. We evaluate three Average Treatment Effect (ATE) estimators: Student’s t-test $\hat{\tau}_{t-test}$, Ordinary Least Squares regression $\hat{\tau}_{OLS}$, and Leave-One-Out Potential outcomes (LOOP) $\hat{\tau}_{LOOP}$. Our approach can be used for future E-TRIALS experiments. Code used for this work is available at https://osf.io/xp6ch/.
Keywords: Educational Interventions, E-TRIALS, A/B Testing, Learning Analytics
@@ 1 Introduction
In 1923, Thomas Edison predicted that motion pictures would revolutionize the education system within 20 years by replacing teachers and textbooks entirely [Salisbury (1986)]. This prediction reflects a recurring pattern in the education system, which is the introduction of new technologies often sparks enthusiasm and bold claims about their potential to transform learning. For instance, in 1961, the Ford Foundation invested over $10$ million in a program to promote televised instruction. The foundation believed televised instruction could ”equal or surpass conventional methods” and make superior teaching resources accessible to more students [Ford Foundation (1961)]. Some of the funds were even used to release teachers from their daily duties to serve as ”television consultants.”
@@
The growing presence of computers in schools paved the way for intelligent tutoring systems (ITS), which aimed to deliver personalized instruction, provide immediate feedback, and assess student knowledge [Billings and Moursund (1988)]. One of the earliest examples, SOPHIE (SOPHisticated Instructional Environment), was developed in 1975 by Brown et al. SOPHIE simulated electronic equipment and allowed students to practice troubleshooting without the risk of damaging real devices [Brown et al. (1974)]. Leveraging techniques from artificial intelligence, cognitive science, and psychology, SOPHIE was able to understand natural language inputs and respond in a supportive tone within two seconds. Additionally, SOPHIE outperformed human observers in identifying faults and was more effective than students at diagnosing errors in their reasoning.
Another notable advancement in ITS was the development of cognitive tutors based on the Adaptive Control of Thought (ACT) theory of learning and problem-solving [Anderson et al. (1995)]. These tutors were guided by eight principles, such as representing student competence as a set of production rules, providing immediate feedback, and minimizing cognitive load. Researchers applied these principles to create tutors for subjects like LISP, geometry, and algebra. The LISP tutor helped students complete exercises 30% faster and achieve 43% higher scores on a post test. The algebra tutor initially didn’t show significant results, but a later version showed significant achievement gains. The geometry tutor helped improve student’s grades by more than one letter grade on a test.
@@
The adoption of CBLPs in classrooms has surged in recent years, largely due to the increase of availability of computers to each student. Before the COVID-19 pandemic, 45% of public schools reported having a computer for every student [Gray and Lewis (2021)]. By the 2022-2023 academic year, this figure had rise to 94% [Institute of Education Sciences (nd)]. Unlike earlier technologies, CBLPs have consistently demonstrated their ability to enhance learning outcomes, particularly when designed to support rather than replace teachers [Anderson (2014), Murphy et al. (2020), Leelawong and Biswas (2008), Toyama (2011)]. These platforms have transformed how students learn, complete assignments, and receive assessments. For example, CBLPs enable students to receive immediate and personalized feedback, while teachers benefit from analytical dashboards that provide insights into student performance. These tools allow educators to quickly identify struggling students and tailor their support accordingly.
@@
Despite their advantages, developing effective CBLPs is a complex challenge due to the vast array of instructional design choices available. As Koedinger et al., noted in Science, there are approximately 300 trillion possible design configurations [Koedinger et al. (2013)]. Researchers are struggling to identify which instructional choices to include and omit. Instructional choices, such as how often should students receive feedback, and what type of feedback is most effective? Should high-performing students receive delayed feedback, while low-performing students receive immediate feedback? Should instruction rely solely on text, or should it incorporate multimedia elements like videos, graphics, or audio? Are open-ended questions more effective than multiple-choice questions? These are just a few of the many decisions researchers and developers must navigate.
@@
The overwhelming number of possible design choices birthed the term ”the rise of the super experiment” [Stamper et al. (2012)]. This concept refers to data-driven educational experiments conducted at the internet level to identify which instructional choices are most effective and to whom. To address this need, tools like UpGrade by Carnegie Learning and TerraCotta within the Canvas Learning Management System have been developed to facilitate such experiments [Fancsali et al. (2022)]. However, many of these tools are either closed-source or lack the flexibility to explore diverse research questions across various aspects of CBLPs.
In this paper, we introduce Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) a free and open tool developed by ASSISTments. E-TRIALS enables researchers to conduct experiments in any subject area. We describe its features, the types of experiments it supports, and how it can address critical research questions in learning sciences and technologies. Additionally, we highlight how E-TRIALS simplifies data analysis by providing consistent and similar structure of the experiment data. To demonstrate this capability, we selected three experiments in E-TRIALS and used the same code to analyze and report findings. To determine the effectiveness of the experiments, we used three Average Treatment Effect (ATE) estimators: Student’s t-test $\hat{\tau}_{t-test}$, Ordinary Least Squares regression $\hat{\tau}_{OLS}$, and Leave-One-Out Potential outcomes (LOOP) $\hat{\tau}_{LOOP}$.
@@ 2 Experimental Environment
ASSISTments is a web-based and free-to-use CBLP designed to assist teachers, rather than replace them [Heffernan and Heffernan (2014)]. The platform allows teachers to assign mathematical problems from open-source curriculum, such as Illustrative Mathematics, or to create their own problem sets. As students work through problem sets, they can request support, which would appear in the form of hints or explanations. ASSISTments is widely used, with over 500,000 students and 30,000 teachers annually [ASSISTments (2025)].
@@ 2.1 E-TRIALS
The Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) is a built-in research functionality within ASSISTments. It enables researchers to conduct experiments in the realm of learning sciences, without any coding expertise. Before the wide spread of CBLP and tools like E-TRIALS, researchers relied on manual methods to design and conduct experiments [Angrist (2003)]. Researchers would rely on a coin toss to assign students to a condition. Fortunately, E-TRIALS simplifies this process of conducting experiments and provides researchers with in-depth data about the student. With over 500,000 active student users and more than 30,000 teacher users [ASSISTments (2025)], ASSISTments offers researchers access to diverse study participants from various background.
@@ 2.1.1 Previous Experiments in E-TRIALS
E-TRIALS is one of five digital platforms funded by Institute of Education Sciences (IES) to develop infrastructure supporting research in the field of learning sciences and technologies [E-TRIALS (2025)]. E-TRIALS has already facilitated impactful studies, such as work by [Walkington et al. (2019)], who used E-TRIALS to explore how language skills can impact student performance in algebra and evaluate their effects on accuracy and response time. In the study, they manipulated six language features: 1) the number of sentences, 2) pronouns, 3) word concreteness, 4) word hypernymy, 5) sentence consistency, and 6) problem topic. The study had 451 students and 2,825 observations. Each observation represented a student’s response to a problem. While the findings showed little evidence that language features significantly impact mathematics word problem-solving performance, the findings did reveal that sentence consistency reduce response time.
Another notable study that used E-TRIALS was work by [Vanacore et al. (2023)]. They evaluated the impact of non-cognitive interventions on student learning and behavior. This study included five experiments: 1) embracing mistakes, 2) inspirational quotes, 3) social comparison, 4) emotion labeling, and 5) confidence judgments. The experiments had 20,472 students. Among the five experiments, only social comparison produced a positive effect on post-test performance.
@@
In a broader analysis, [Prihar et al. (2022)] evaluated 50 different experiments conducted using E-TRIALS. These experiments investigated a wide range of research questions, such as the effectiveness of explanation vs. answer-only feedback, video examples vs. text-based, common wrong answer feedback vs. no feedback, fill-in vs. multiple-choice questions, and motivational vs. non-motivation messages. Collectively, these experiments had a 30,408 students. The primary goal of this meta-analysis was to identify consistent patterns across experiments, such as personalization and effective tutoring methods within CBLPs.
@@
These examples are an indication of the types of researcher questions E-TRIALS can answer, and the rich data that researchers receive. E-TRIALS opens new avenues for addressing novel research questions and generating data-driven insights. Its ability to handle large sample sizes, diverse interventions, and complex experimental designs makes it an invaluable tool for advancing research.
@@ 2.1.2 Developing an Experiment
Experiments in E-TRIALS are delivered through ASSISTments. They can be assigned using learning management systems, such as Google Classroom or Canvas. As show in Figure 1, E-TRIALS supports three types of experiments: single support, support comparison, and problem varied.
@@
Single Support: Single support studies are designed to evaluate the effects of a specific type of student support message, such as a hint or explanation. These messages can be in the form of text, images, videos, or audio. The control condition in these studies is the best-so-far condition, which is ASSISTments’ best selection of a support to deliver to maximize the student’s performance.
Researchers use ASSISTments’ math problem content library, which comprises problems from open-source curricula, such as Kendall Hunt, Illustrative Mathematics, and Engage NY/Eureka Math to create experimental conditions. Researchers do not need to recruit students because they’re randomly assigned to a condition.
Support Comparison: Support comparison studies build on the single support model by allowing researchers to investigate two or more types of student support messages. Examples include comparing hints vs. explanations, hints vs. worked examples, or explanations vs. a dynamic algebraic notation system. As with single support studies, researchers can use ASSISTments’ existing math content to create conditions. For example, [Smith et al. (2022)] used a support comparison study to explore the impact of worked example extensiveness and the degree of dynamic presentation on student learning. Their study had six conditions: 1) static concise, 2) static extended, 3) sequential concise, 4) sequential extended, 5) dynamic history, and 6) dynamic no history. The results revealed an overall improvement from pretest to post test after viewing the worked examples, but there was no significant differences on post test performance between worked example presentation.
Problem Varied: Problem varied studies offer researchers the flexibility to create their own problem content and recruit participants. These studies accommodate a broader range of participants, from K-12 to higher education students. [Smalenberger and Smalenberger (2022)] conducted a problem varied study to examine how middle and college students benefited from three types of problem support messages: text hints, explanations, and video explanations. They compared the effects of these formats within the college sample, and between the college and middle school samples. Findings revealed that for college students, on-demand video explanations was more effective than text hints. Problem varied studies are not limited to math-related research questions. For example, [Unal et al. (2020)] used this study type to design a think-aloud study that explored cognitive control and rule learning in a real-world setting. Their study featured probability problem sets and recruited undergraduate students as participants.

### v2

@@ Abstract
Computer-based learning platforms (CBLPs) have become a common medium in schools, transforming how students learn and interact with educational content. However, researchers still lack adequate tools to address the diverse set of challenges that students face in these environments. In this paper, we introduce Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS), a free tool developed by ASSISTments to help researchers conduct randomized controlled trials in the realm of learning sciences. We describe its features, the types of experiments it supports, and how it can address critical research questions. We showcase E-TRIALS’ capabilities through two real-world interventions. Finally, we evaluate the efficacy of interventions using three average treatment effect (ATE) estimators. Student’s t-test, regression, and Leave-One-Out Potential outcomes (LOOP). The results demonstrate that the unbiased LOOP estimator can achieve greater precision by adjusting for baseline covariates compared to the Student’s t test. Our work demonstrates the potential of E-TRIALS to advance research and contribute to the development of more effective, inclusive, and adaptive CBLP. The code used for this work is available at https://osf.io/xp6ch/.
Keywords: Educational Interventions, E-TRIALS, A/B Testing, Learning Analytics, Treatment Effect Estimators
@@ 1 Introduction
In 1923, Thomas Edison predicted that motion pictures would revolutionize the education system within 20 years by replacing teachers and textbooks entirely [Salisbury (1986)]. This prediction reflects a recurring pattern in the education system, which is the introduction of new technologies often sparks enthusiasm and bold claims about their potential to transform learning. For instance, in 1961, the Ford Foundation invested over $10$ million US dollars - equivalent to approximately $106$ million US dollars in 2025 - in a program aimed to promote televised instruction. The foundation believed televised instruction could ”equal or surpass conventional methods” and make superior teaching resources accessible to more students [Ford Foundation (1961)]. Some of the funds were even used to release teachers from their daily duties to serve as ”television consultants.”
@@
The growing presence of computers in schools paved the way for intelligent tutoring systems (ITS), which aimed to deliver personalized instruction, provide immediate feedback, and assess student knowledge [Billings and Moursund (1988)]. One of the earliest examples, SOPHIE (SOPHisticated Instructional Environment), was developed in 1975 [Brown et al. (1974)]. SOPHIE simulated electronic equipment and allowed students to practice troubleshooting without the risk of damaging real devices [Brown et al. (1974)]. Leveraging techniques from artificial intelligence, cognitive science, and psychology, SOPHIE was able to understand natural language inputs and respond in a supportive tone within two seconds. Additionally, SOPHIE outperformed human observers in identifying faults in their trouble-shooting reasoning and was more effective than students at diagnosing errors in their reasoning.
Another notable advancement in ITS was the development of cognitive tutors based on the Adaptive Control of Thought (ACT) theory of learning and problem-solving [Anderson et al. (1995)]. These tutors were guided by eight principles, such as representing student competence as a set of production rules, providing immediate feedback, and minimizing cognitive load [Anderson et al. (1995)]. Researchers applied these principles to create tutors for subjects like LISP, geometry, and algebra [Anderson et al. (1995)]. The LISP tutor helped students complete exercises 30% faster and achieve 43% higher scores on a posttest [Anderson et al. (1995)]. The algebra tutor initially didn’t show significant results, but a later version showed significant achievement gains. The geometry tutor helped improve student’s grades by more than one letter grade on a test [Anderson et al. (1995)].
@@
The adoption of CBLPs in classrooms has surged in recent years, largely due to the increase of availability of computers to each student [Gopika and Rekha (2023)]. Before the COVID-19 pandemic, 45% of public schools reported having a computer for every student [Gray and Lewis (2021)]. By the 2022-2023 academic year, this figure had risen to 94% [Institute of Education Sciences (nd)]. Unlike earlier technologies, CBLPs have consistently demonstrated their ability to enhance learning outcomes, particularly when designed to support, rather than replace teachers [Anderson (2014), Murphy et al. (2020), Leelawong and Biswas (2008), Toyama (2011)]. CBLPs have transformed how students learn, complete assignments, and receive assessments. For example, CBLPs enable students to receive immediate and personalized feedback, while teachers benefit from analytical dashboards that provide insights into student performance. These tools allow educators to quickly identify struggling students and tailor their support accordingly.
@@
Despite their advantages, developing effective CBLPs is a complex challenge due to the vast array of instructional design choices available. As Koedinger et al., noted in Science, there are approximately 300 trillion possible design choices [Koedinger et al. (2013)]. Researchers are struggling to identify which instructional choices to include and omit. Instructional choices, such as how often should students receive feedback, and what type of feedback is most effective? Should high-performing students receive delayed feedback, while low-performing students receive immediate feedback? Should instruction rely solely on text, or should it incorporate multimedia elements like videos, graphics, or audio? Are open-ended questions more effective than multiple-choice questions? These are just a few of the many decisions researchers and developers must navigate.
@@
The overwhelming number of possible design choices birthed the term ”the rise of the super experiment” [Stamper et al. (2012)]. This concept refers to data-driven educational experiments conducted at the internet level to identify which instructional choices are most effective and to whom. To address this need, tools like UpGrade by Carnegie Learning and TerraCotta within the Canvas Learning Management System have been developed to facilitate such experiments [Fancsali et al. (2022)]. However, TerraCotta is used for experiments within a learning management system like as Canvas or Google Classroom [Motz et al. (2024)], and UpGrade is a tool that works by using its API to connect to the CBLP to deliver the educational content and researchers would need to develop their own content [Wang et al. (2019)].
In this paper, we introduce Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) a free and open tool developed by ASSISTments. E-TRIALS enables researchers to conduct experiments in any subject area. We describe its features, the types of experiments it supports, and how it can address critical research questions in learning sciences and technologies (LS&T). To demonstrate the capability of E-TRIALS, we introduce two E-TRIALS experiments and evaluate the effectiveness of the interventions using three Average Treatment Effect (ATE) estimators: (1) Student’s t-test ($\hat{\tau}_{\text{t-test}}$), (2) Ordinary Least Squares regression ($\hat{\tau}_{\text{Reg}}$), and (3) Leave-One-Out Potential outcomes estimator (LOOP) ($\hat{\tau}_{\text{LOOP}}$). These three methods are chosen because they represent a progression in complexity and robustness. The t-test estimator is the simplest and most intuitive, relying solely on the difference in means between the treatment and control groups. Although RCTs should, in theory, ensure covariate balance on average, real-world RCTs often exhibit residual imbalance. In such cases, covariate adjustment becomes important. Regression is the most commonly used method for this purpose. The LOOP estimator, a more advanced approach, offers an unbiased estimate of the ATE while accounting for covariate differences, even in small samples.
@@ 2 Experimental Environment
ASSISTments is a web-based and free-to-use CBLP designed to assist teachers, rather than replace them [Heffernan and Heffernan (2014)]. The platform allows teachers to assign mathematical problems from open-source curriculum, such as Illustrative Mathematics, or to create their own problem sets. As students work through their assigned assignments, they can request support, which would appear in the form of hints or explanations. ASSISTments is widely used, with over 500,000 students and 30,000 teachers annually [ASSISTments (2025)].
@@ 2.1 E-TRIALS
The Ed-Tech Research Infrastructure to Advance Learning Sciences (E-TRIALS) is a research tool designed by ASSISTments. It enables researchers to conduct experiments in the realm of learning sciences in any subject without writing a single line of code. Before the wide spread of CBLP and tools like E-TRIALS, researchers relied on manual methods to design and conduct experiments [Angrist (2003)]. E-TRIALS connects researchers with a large and diverse population of students (¿500,000) and teachers (¿30,000) to conduct experiments. Once a researcher completes designing the experiment in E-TRIALS, they would not need to worry about recruiting participants because ASSISTments does the recruiting for them by assigning its users to the experiment conditions. E-TRIALS is one of five digital platforms funded by the Institute of Education Sciences (IES) to develop an infrastructure to support educational research questions [Schellinger et al. (2024)]. Altogether, these platforms - UpGrade, Kinetic, Terracotta, and ASU’s Learning@Scale - have helped researchers connect to millions of teachers and learners across K-16 [Schellinger et al. (2024)].
@@ 2.1.1 Previous Experiments in E-TRIALS
Since the inception of E-TRIALS, dozens of researchers have used it to conduct RCTs [E-TRIALS (2025)]. Walkington et al. conducted a study in E-TRIALS that explored how language skills can impact student performance in algebra and evaluate their effects on accuracy and response time [Walkington et al. (2019)]. They manipulated six language features: 1) the number of sentences, 2) pronouns, 3) word concreteness, 4) word hypernymy, 5) sentence consistency, and 6) problem topic. The study had 451 students and 2,825 observations, where each observation represented a student’s response to a problem. While the findings showed little evidence that language features significantly impact mathematics word problem-solving performance, the findings did reveal that sentence consistency reduced response time. Another study on algebra explored the effects of feedback for middle school students by testing four conditions across two experiments: 1) students received correct-answer feedback or no feedback, and 2) students received explanation feedback or try-again feedback [Fyfe (2016)]. Across a total of 246 students in both experiments, posttest scores revealed that students with low prior knowledge benefited from all feedback types, instead of the no feedback condition. However, all four conditions didn’t affect the performance of students with high prior knowledge. Though these two studies demonstrate E-TRIALS’ ability to help researchers improve student performance through research-backed interventions, the following study is especially worth mentioning due to the large number of student participants of 2,152 from 5th-12th grades [Harrison et al. (2020)]. Students were randomly assigned to one of four conditions that manipulated the spacing between numbers and terms within mathematical expressions to be either 1) neutral (no extra spaces), 2) congruent with the order of operations (using grouping), 3) incongruent with the order of operations, or 4) a combination of the previous conditions. By using mastery speed (students’ total attempt count before correctly answering three consecutive problems) as the outcome measure, the study found that students in the incongruent condition had slower mastery speed than those in the congruent or neutral conditions. These findings suggest that spacing that conflicts with the order of operations presents additional challenges for students when solving mathematical expressions.
In the above papers, researchers did not have to recruit teachers or students, or even design the problem content. They selected a set of problem set(s) from the ASSISTments library of math content and then incorporated their condition(s) into the problem set(s). Once the study is ready, ASSISTments deploys it and randomizes the students to a condition. E-TRIALS also offers the flexibility for researchers to create their own problem content and recruit participants. A study by [Oppenzato (2024)], noticed that oftentimes addition and multiplication fraction problems in textbooks are often common for some combinations of operand types and operations, while it can be rare for others. In the first iteration of the study, the researcher designed a pretest-intervention-posttest in Qualtrics to find out whether a more balanced distribution of practice problems, especially for rare types of problems could improve student performance. 63 students were randomly assigned to one of two conditions: 1) more common types of problems, or 2) balanced mix common and rare problems. The final data contained 40 students because many were removed for scoring a perfect score on the pretest. Due to a small sample size, the researcher decided to redo the study. The second iteration of the study was conducted in E-TRIALS to ensure a larger sample size and only allow teachers who have a high rate of underperforming students to participate in the study. The researcher developed the content for the pre- and posttest, and the practice problems. Findings from 127 students show that students who received the balanced mix of common and rare problems improved and made fewer independent whole number errors on whole-fraction multiplication items than students who received common types of problems.
Another study by Zhou et al. demonstrates another advantage E-TRIALS offers researchers [Zhou et al. (2021)]. Zhou et al. investigated how hint and scaffolding tutoring strategies can influence adult learners’ interactions within the edX MOOC platform. Using the Learning Tools Interoperability, edX and ASSISTments were integrated to deliver the content via edX and the assignment via ASSISTments. Students in the first condition were able to request hints at any point during a problem. Students in the second condition automatically received a sequence of scaffolding questions if their first attempt on an incorrect problem, and they were required to complete the entire sequence before starting the next problem. Results from the study revealed that scaffolding didn’t help adult students. In fact, students with low prior knowledge requested significantly more correct answers and failed to solve later problems.
@@
In a broader analysis, [Prihar et al. (2022)] aggregated results from 50 experiments that were conducted in E-TRIALS from various researchers. These experiments investigated a wide range of research questions, such as the effectiveness of explanation vs. answer-only feedback, video examples vs. text-based, common wrong answer feedback vs. no feedback, fill-in vs. multiple-choice questions, and motivational vs. non-motivation messages. Collectively, these experiments had 30,408 K-12 students. The primary goal of this meta-analysis was to identify consistent effective patterns across experiments, such as personalization and effective tutoring methods within CBLPs.
@@
As demonstrated by the interventions discussed above, E-TRIALS has enabled researchers to explore researcher questions with minimal effort. Its flexibility coupled with the types of data the platform collects has allowed researchers to offer researched-backed design choices for CBLP, and in doing so, helped learners perform better academically. E-TRIALS opens new avenues for addressing novel research questions and generating data-driven insights. Its ability to handle large sample sizes, diverse interventions, and complex experimental designs makes it an invaluable tool for advancing research.
@@ 2.1.2 Developing an Experiment
Experiments in E-TRIALS are delivered through ASSISTments. They can be assigned using learning management systems, such as Google Classroom or Canvas. As shown in Figure 1, E-TRIALS supports three types of experiments: single support, support comparison, and problem varied.
@@
Single Support: Single support studies are designed to evaluate the effects of a specific type of student support message, such as a hint or explanation. These messages can be in the form of text, images, videos, or audio. The control condition in these studies is the best-so-far condition. This condition represents the highest-performing student support currently in ASSISTments for that problem.
Researchers use ASSISTments’ math problem content library, which comprises problems from open-source curricula, such as Kendall Hunt, Illustrative Mathematics, and Engage NY/Eureka Math to create experimental conditions [Illustrative Mathematics (2025), Great Minds (2025)]. Once they create the experimental conditions in a problem set(s), teachers can assign the problem set(s) to their students, and in doing so, students would randomly be assigned to a condition. This process alleviates the burden of participant recruitment.
Support Comparison: Support comparison studies build on the single support model by allowing researchers to investigate two or more types of student support messages. Examples include comparing hints vs. explanations, hints vs. worked examples, or explanations vs. a dynamic algebraic notation system. As with single support studies, researchers can use ASSISTments’ existing math content to create conditions. For example, [Smith et al. (2022)] used a support comparison study to explore the impact of worked example extensiveness and the degree of dynamic presentation on student learning in algebra. Their study had six conditions: 1) static concise, 2) static extended, 3) sequential concise, 4) sequential extended, 5) dynamic history, and 6) dynamic no history. Data from 230 7th-12th graders revealed an overall improvement from the pretest to the posttest after viewing the worked examples, but there were no significant differences in posttest performance between worked example presentations. Meaning, the extensiveness and the degree of dynamic presentation did not impact students’ posttest scores.
Problem Varied: Problem varied studies offer researchers the flexibility to create their own problem content and recruit participants. These studies accommodate a broader range of participants, from K-12 to higher education students. This option is suitable for researchers that want to explore research questions outside the realm of mathematics, K-12 students, or would like to use a platform other than ASSISTments to deliver the assignment. [Smalenberger and Smalenberger (2022)] conducted a problem varied study to investigate whether findings from middle school students can be replicated among collect level students. Both middle and college-level students were assigned to one of three tutor strategies conditions: 1) text hints, 2) text explanations, or 3) video explanations. Findings revealed that college students benefited from video explanations more than text hints. However, middle school students benefited more from video explanations than college students by 36%. [Unal et al. (2020)] also used a problem varied study to design a think-aloud intervention that explored cognitive control and rule learning in a real-world setting. Their study featured probability problem sets and recruited undergraduate students as participants.